A balloon starts riding from ground from rest with an upward acceleration 2 m/s^2. Just after 1 s, a stone is dropped from it. The time taken be stone to strike the ground is nearly1) 0.3 s2) 0.7 s3) 1 s4) 1.4 sWhy we didnt used the value of acceleration as 2 m/s^2 or 2+10 to calculate s?

Grade:12th pass

1 Answers

Arun
5 years ago
for balloon,
initial velocity of balloon=0  so,u=o
a=2m/s2
t=1s
v=u+at                
v=2m/s  (as u=0m/s ,a=2m/s2,t=1s)
s=ut+1/2at2      
s=1m     (as u=0m/s ,a=2m/s2,t=1s)
for stone,
u=-2m/s ,s=1m,a=9.8m/s2
at2+2ut-2s=0
9.8t2-4t-2=0
t=.7s